### v3

# Yushi Hu*1 2] Weijia Shi1Xingyu Fu3Dan Roth3Mari Ostendorf1 Luke Zettlemoyer1Noah A. Smith*1 2] Ranjay Krishna*1 2]

Abstract
Humans draw to facilitate reasoning: we draw auxiliary lines when solving geometry problems; we mark and circle when reasoning on maps; we use sketches to amplify our ideas and relieve our limited-capacity working memory. However, such actions are missing in current multimodal language models (LMs). Current chain-of-thought and tool-use paradigms only use text as intermediate reasoning steps. In this work, we introduce Sketchpad, a framework that gives multimodal LMs a visual sketchpad and tools to draw on the sketchpad. The LM conducts planning and reasoning according to the visual artifacts it has drawn. Different from prior work, which uses text-to-image models to enable LMs to draw, Sketchpad enables LMs to draw with lines, boxes, marks, etc., which is closer to human sketching and better facilitates reasoning. Sketchpad can also use specialist vision models during the sketching process (e.g., draw bounding boxes with object detection models, draw masks with segmentation models), to further enhance visual perception and reasoning. We experiment on a wide range of math tasks (including geometry, functions, graph, chess) and complex visual reasoning tasks. Sketchpad substantially improves performance on all tasks over strong base models with no sketching, yielding an average gain of 12.7% on math tasks, and 8.6% on vision tasks. GPT-4o with Sketchpad sets a new state of the art on all tasks, including $V^{*}$ench (80.3%), BLINK spatial reasoning (83.9%), and visual correspondence (80.8%).All codes and data are in https://visualsketchpad.github.io/.
1 Introduction
Sketching is a fundamental human activity, serving as a versatile tool for communication [11], ideation [48], and problem-solving [47]. Unlike written language, sketches have the advantage of conveying visuo-spatial ideas directly, for example by using spatial relations on paper to convey spatial relations or other more abstract relationships in the world. This may explain their ubiquity; maps [40] and architectural plans [12] have been found incised in stone, etched on leather, impressed in clay, and drawn on paper in diverse cultures scattered across the world [39]. Sketches are so fundamental that we use them to teach school children how to solve geometry problems by drawing support lines, to aid engineers conveying prototypes, to support architects creating blueprints, and to allow scientists like us to convey scientific contributions (see Figure 1).
Figure 1: Sketchpad equips GPT-4 with the ability to generate intermediate sketches to reason over tasks. Given a visual input and query, such as proving the angles of a triangle equal 180°, Sketchpad enables the model to draw auxiliary lines which help solve the geometry problem. The examples are from [8, 51, 44]. For all these examples, without Sketchpad, GPT-4o fails to get the correct answer, while Sketchpad + GPT-4o achieves the correct solution.
As multimodal language models (LMs) [35, 42, 28, 27, 1, 3, 49, 7, 31, 41, 6, 5] have begun to mature, we now expect them to solve tasks like the ones mentioned above, i.e., ones where people draw intermediate sketches to simplify reasoning. Popular benchmarks now include questions about geometry (e.g., Geometry3K [32]) and complex math problems (e.g., IsoBench [8]). In these benchmarks, models are given images of diagrams and asked questions requiring symbolic grounding and spatial understanding, where intermediate sketches like auxiliary lines can enhance reasoning. Even benchmarks in computer vision now have a similar flavor. Specialist vision models can be viewed as sketching on natural images. For example, object detection is plotting bounding boxes around objects; depth estimation is drawing colormaps according to depth. The recently proposed BLINK benchmark [9] would benefit significantly from such intermediate visual sketches. Similarly, the V*Bench benchmark [51] could focus reasoning on image crops to find answers. Unfortunately, current LMs lack a scaffold for using sketch-based reasoning when solving tasks.
In this paper, we introduce Visual Sketchpad: a framework that provides multimodal LMs with the tools necessary to generate intermediate sketches to reason over tasks. Inspired by textual chain-of-thought reasoning in LMs [50, 62], Sketchpad prompts the underlying visual LM to produce visual artifacts as part of a chain of mixed textual, programmatic, and visual reasoning. For example, to prove that the angles of triangles sum up to 180 degrees in Figure 1 (a), Sketchpad enables agents to modify the diagram by introducing a new auxiliary line. This new line, along with new annotated angles, provides the critical information to solve the geometry task. Similarly, Sketchpad improves models’ spatial reasoning for computer vision. To determine if there are cookies stacked on top of other cookies in the image (Figure 1b), the model first produces an intermediate depth estimate. By analyzing the depth estimate, which reveals cookies overlapping at different depths, the model is able to correctly answer that the cookies are indeed stacked.
We demonstrate the effectiveness of visual Sketchpad across a wide range of mathematics and computer vision tasks. For math, we tackle problems including (1) geometry  [32], (2) mathematical functions, (3) graph algorithms, and (4) strategy games  [8]. For geometry questions, Sketchpad enables models to generate Matplotlib code with auxiliary lines and variables, given the diagram input and questions (Figure 1a). Notably, even when the input is solely language-based, such as mathematical functions, Sketchpad enables models to plot the functions and reason about their properties, using only the mathematical function expression as input (Figure 1b). These results highlight the ability of Sketchpad to aid reasoning, even in tasks with purely language-based inputs. Across all four categories of mathematical tasks, Sketchpad consistently improves the baseline GPT-4o performance, yielding an average gain of 11.2%. For computer vision, we tackle diverse tasks including (1) depth, (2) spatial reasoning, (3) jigsaw, (4) visual correspondence, (5) semantic correspondence, as well as questions from (6) the MMVP and (7) the V*Bench benchmarks [9, 44, 51]. For this domain, Sketchpad enables models to generate segmentation masks, crop images, draw bounding boxes, zoom into image regions, overlay images, etc. Similar to math, Sketchpad shows consistent improvements across all seven types of computer vision tasks. For example, GPT-4o, augmented with Sketchpad, sees 14.3% improvement on V*Bench, 12.1%, and 9.7% improvements on BLINK’s depth and semantic correspondence tasks, setting a new state of the arts across all tasks. Finally, we analyze the effectiveness of Sketchpad by comparing the plans generated by our model with human-created plans, showing that they are well-aligned and exhibit similar reasoning patterns. We hope Sketchpad opens up new research opportunities toward more capable and interpretable multimodal intelligence.
2 Related Work
Sketchpad generalizes recent work on multimodal tool-use and visual prompting. We also place our work within the larger sphere exploring LMs as agents.
Visual programming and tool-use. With the advancement of LMs [4, 35, 42, 45, 13], researchers have demonstrated the possiblity of decomposing complex vision tasks into simpler substeps that can each be solved using vision tools [57, 60, 18, 17]. Among them, the most relevant to us are Visprog [14] and ViperGPT [38]. They use LMs to generate Python code, which sequencially invokes specialized vision tools. These methods share a common problem that the multimodal modules follow a pre-defined plan outlined by the LM. By contrast, Sketchpad allows LMs to change their plan according to the intermediate visual artifacts, yielding better performance and robustness when solving complex multimodal tasks.
Visual prompting. Recent work shows that multimodal models can be augmented by visual prompts added to natural images [37]. For example, SoM [55] shows that adding labeled segmentation masks on images unleashes GPT-4V’s visual grounding ability. Prior work also reports similar findings in 3D [25] and Robotics [34]. Sketchpad is a generalized framework for all these methods, allowing LMs to decide what visual prompting to use as part of the multimodal reasoning process.
LMs as agents. Recent work has started to treat LMs as agents that can both reason and act [58, 33, 52, 36, 59]. Researchers have applied this idea to software engineering [20, 61, 15], robotics [34], vision [29, 57], and GUI navigation [54, 23, 53]. Sketchpad can also be viewed as an agent that accepts multimodal inputs and outputs. One big difference is that Sketchpad can create visual artifacts to facilitate reasoning, while prior LM agents only generate text during reasoning.
3 Visual Sketchpad
Figure 2: Overview of Sketchpad. Given a multimodal query, the Sketchpad agent generates a sketching plan to address the query (Thought), and then synthesizes a program to create visual sketches (Action). By analyzing the resulting sketches (Observation), which serve as a visual representation of the reasoning process, the model generates a final response to the query.
We introduce visual Sketchpad, a general framework that enables multimodal LMs to draw sketches as intermediate reasoning steps and to use these sketches to facilitate further reasoning. Figure 2 shows examples of how Sketchpad works. Given a multimodal query, Sketchpad agent generates a sketching plan to address the query (Thought), and then synthesizes a program to create visual sketches (Action). By analyzing the resulting sketches (Observation), which serve as a visual representation of the reasoning process, the model generates a final response to the query.
Our framework requires no finetuning or training. Multimodal LMs, out of the box, can be prompted to sketch using our framework. Our implementation is based on the AutoGen [52] framework. We give the overview of our Sketchpad framework in §3.1, and delve deep into how it integrates sketching into the reasoning process in §3.2.
3.1 Overview of Sketchpad
The Sketchpad agent solves tasks by engaging in an iterative interaction process with an environment. Given a multimodal query $q$ that includes both visual and textual components, the model generates a series of thoughts, actions, and observations to gather the information needed to answer the query. At each time step $t$, the model performs three key steps:
Thought: The model analyzes the current context $c_{t}$, which includes the query, previous thoughts, actions, and observations, to generate a thought plan $p_{t}$ for the next action. For example, given the query $q$ – “find the $\angle EIC$” in Figure 2a, the model’s thought plan $p_{1}$ is to draw an auxiliary line $IX$ parallel to $BD$ serving as a visual sketch to help solve the problem.
Action: Based on the thought plan, the model executes an action $a_{t}$, which can manipulate both visual and textual content. In the geometry example, to realize the proposed thought of drawing the auxiliary line, the model generates Python code to modify the original geometry diagram. The generated code is then compiled and executed.
Observation: Based on the action $a_{t}$, Sketchpad’s environment returns a new observation $o_{t+1}$, such as a new diagram with the auxiliary line drawn in the geometry example. The multimodal context is then updated to $c_{t+1}=(c_{t},p_{t},a_{t},o_{t+1})$.
The multi-turn interaction process continues until time step $T$, when the model determines that it has gathered enough information from the context $c_{T}$ to answer the query. At this point, it generates a special Terminate action and provides the answer.
Different from prior work [58], where LMs primarily generate and manipulate text-based observations and actions, Sketchpad enables the model to work with multimodal observations $o_{t}$ and actions $a_{t}$, manipulating both visual and textual content. This allows the model to plan and reason with the visual sketches it has drawn, enhancing its problem-solving capabilities.
3.2 Sketching via Code Generation
The core component of Sketchpad is sketching, which enables the LM to generate visual sketches by synthesizing programs that call different specialist vision models or Python plotting packages.
Program Generation. Similar to recent works like ViperGPT and VPD [14, 38, 18], Sketchpad enables LMs to sketch through code generation. The LM is provided, through a prompt, with a detailed description of the available tools that can generate multimodal content (an example prompt and description can be found in §C). The prompt includes Python function signatures and docstrings [16] for these modules, but does not contain their full implementation. The LM generates Python code in a code block, using the provided tools, which, when executed, generates new image and text outputs. A special display function allows the LM to visualize the sketch image in the next observation $o_{t+1}$.
Modules for sketching. Sketchpad uses various tools to facilitate the sketching process, depending on the task at hand. For mathematical tasks, Sketchpad uses common Python packages like matplotlib and networkx for plotting (see §4). For vision tasks, the LM leverages specialist vision models during the sketching process. These models include detection tools that draw bounding boxes on the image, as well as segmentation and marking tools (inspired by SoM [55]) that draw colorful masks on the image and use numbers to label segments. We find these specialists possess useful perception skills for visual reasoning tasks, and Sketchpad is an effective way to combine them into a multimodal LM (see §5.1).
4 Sketching to Solve Math Problems
In this section, we experiment with Sketchpad on four complex mathematical tasks : (1) geometry, (2) mathematical functions, (3) graph algorithms, and (4) game strategies. We demonstrate that incorporating sketching capabilities into LMs significantly improves their performance on these mathematical problems, setting new state-of-the-art results (§4.1).
Details of our evaluation tasks and the tools employed for visual reasoning are as follows:
Geometry Problems. Drawing auxiliary lines in geometry diagrams is often helpful for problem-solving. For example, in Figure 2 (a), when asked to find $\angle EIC$, the LM plans to draw an auxiliary line $IX$ parallel to $BD$, allowing it to use the properties of parallel lines to determine $\angle EIC$. To evaluate the effectiveness of Sketchpad, we use the problems from the Geometry3K dataset [32].
To realize the line drawing process, Sketchpad takes a geometry diagram and its corresponding matplotlib code as input. The model then proposes and modifies the code to generate auxiliary lines, and executes the modified code to visualize the updated diagram with the added lines.
Mathematical functions. Understanding mathematical functions is crucial for various applications in science, engineering, and economics. We focus on two tasks related to mathematical functions from the IsoBench datasets [8]:
  • Classifying parity aims to determine whether a function is even, odd, or neither. Even functions satisfy $f(-x)=f(x)$ for all $x$, while odd functions satisfy $f(-x)=-f(x)$.
  • Identifying convexity/concavity aims to determine whether a function is convex or concave.
Existing LMs can only analyze functions and attempt to prove their properties analytically.  [^1] For humans, the analytical approach is the correct way to tackle these math tasks. However, we observe that LMs are not good at analytical reasoning in math. They make errors when deducing $f(-x)$ and derivatives. However, Sketchpad enables them to visually sketch functions to solve problems more efficiently. For instance, to determine the convexity of the function in Figure 1b, Sketchpad allows the model to plot the function using matplotlib, and visually inspect its overall shape.
Graph algorithms. Many real-world problems, such as those related to computer networks and transportation systems, can be formulated as graph problems. We evaluate Sketchpad on three graph problems from IsoBench [8]:
  • Graph connectivity determines whether there exists a path between two vertices in a graph.
  • Maximum flow aims to find the maximum amount of flow that can be sent through a network from a source vertex to a sink vertex, subject to capacity constraints on the edges.
  • Graph isomorphism tests whether two graphs are structurally equivalent.
Given an adjacency matrix of a graph like in Figure 2(b), Sketchpad can draw the actual graph structure, using using Python’s networkx library, enabling direct visual reasoning about graph properties and relationships.
Game strategies. Chess games can be represented in various formats, including visual board states and textual move notations. Given only the textual move notations, Sketchpad can draw the visual representations of the chess board to analyze positions and formulate strategies. We evaluate the performance of Sketchpad on the winnder identification task from the IsoBench datasets [8] that aims to find the outcome of a chess game (win for White, win for Black, or draw) based on the final board state. To create the graphical board, Sketchpad uses Python’s chess library to draw the board from the Forsyth-Edwards Notation (FEN) of chess.
Geometry Graph Math Game
Model Geometry Maxflow Isomorphism Connectivity Convexity Parity Winner ID
Prior LLMs without visual inputs
[…]
Latest multimodal LLMs + Visual Sketchpad
GPT-4 Turbo 37.5 32.8 62.5 66.0 57.0 80.5 50.4
+ Sketchpad 45.8 96.8 97.6 97.6 77.3 71.5 64.2
+8.3 +64.0 +35.1 +31.6 +20.3 -9.0 +13.8
GPT-4o 62.5 25.0 50.8 96.1 87.2 84.4 61.1
+ Sketchpad 66.7 66.3 65.3 98.4 94.9 94.7 64.6
+4.2 +41.3 +14.5 +2.3 +7.7 +10.3 +3.5
Table 1: Accuracy scores on geometry problems, graph algorithms, mathematical functions, and game. Sketchpad yields large performance gains on most tasks and outperform all baselines.
4.1 Results
We evaluate the performance of Sketchpad on multimodal LMs with API access, including gpt-4-turbo-2024-04-29 and gpt-4o-2024-05-13. We compare these results to baselines without the Visual Sketchpad and other notable closed-source models, such as Claude 3 and Gemini-Pro, as well as open-source models like Mistral [19] and LLaMA-2 70B [46].
Main results. As shown in Table 1, Sketchpad consistently improves base model performance across all tasks, with an average improvement of 11.2% for GPT-4o and 23.4% for GPT-4 Turbo. In particular, we observe large gains on graph algorithms such as maximum flow and connectivity. For instance, GPT-4o with Sketchpad achieves an accuracy of 66.3% on the maximum flow problem, improving over the base model by 41.3%. Similarly, Sketchpad substantially improves the performance on mathematical functions, with GPT-4 Turbo achieving over 90% accuracy and GPT-4o over 88% accuracy on convexity and parity classification tasks. Furthermore, we observe gains (3% $\sim$ 10%) on game strategies, demonstrating that drawn game boards drawn can improve reasoning about game strategies. Overall, these results highlight the effectiveness of Sketchpad in enhancing the reasoning capabilities of multimodal language models across diverse domains.
5 Sketching to Solve Computer Vision Tasks
In this section, we experiment with Sketchpad on complex visual reasoning tasks. Recent work (BLINK) [9] finds that many core visual perception abilities are still missing from existing multimodal LMs—even though many computer vision specialist models possess such abilities. Also, SoM [55] shows that drawing segmentation masks on images unleashes the strong visual grounding ability of GPT-4V. We generalize these ideas with Sketchpad, allowing LMs to use specialist vision models to sketch. Details of these modules are in §5.1. Sketchpad enhances multimodal LMs’ visual reasoning abilities and establishes new SOTAs on all 7 tasks (§5.2).
Tasks. We experiment with a wide range of complex visual reasoning tasks: (1) $V^{*}$ench [51]. This benchmark contains questions about small items in an image. (2) MMVP benchmark from Eyes Wide Shut [44]. This benchmark contains visual questions specially designed to reveal the visual shortcomings of CLIP-based multimodal LMs. (3) BLINK [9]. This benchmark contains visual perception tasks that are easy for humans, but post significant challenge for multimodal LMs. Specifically, we experiment with relative depth, spatial reasoning, jigsaw puzzle, visual correspondence, and semantic correspondence tasks. More details of each task are in §D.
Figure 3: Examples of Sketchpad applied to vision tasks. The figure shows actual outputs generated by Sketchpad. By contrast, the baseline GPT-4o model cannot answer these questions correctly. Note that for demonstration purposes, the “A” and “B” marks in (a) are different from the actual images in the experiments.
5.1 Vision Specialists as Sketching Tools in Sketchpad
LMs can use the following modules to sketch and manipulate images. We wrap these modules into Python functions that the LMs can call. Refer to §C for the function definitions.
Detection. This module takes an image and a simple text query (e.g., “cat”) as input. We run the Grounding-DINO [30] open-vocabulary objection detection model and plot the detected bounding boxes (together with a number label) on the image. It also returns the bounding box coordinates.
Segmentation. This module takes an image as input and returns an image with colorful segmentation masks on it. Each mask also has a number label. We follow the implementation of SoM [55]. The underlying segmentation models are SegmentAnything [22] and Semantic-SAM [24].
Depth estimation. This module takes an image as input and returns a depth map. The underlying model is DepthAnything [56].
Visual search via sliding window. This module mimics how humans search for small items on an image. It takes a text query as input and runs a sliding window over the image. The window size is 1/3 of the image size, and the step size is 2/9 of the image size (so an image will have $4\times 4=16$ windows). It returns a sequence of image patches in which the query is detected.
Other image manipulation modules. Other modules include (1) zoom-in and crop, which takes an image and a bounding box as input and returns the image patch inside the box; (2) Overlay images, which takes two images and alpha values as input, and returns the overlayed image.
5.2 Results
We experiment with the same multimodal LMs as in §4 on complex visual reasoning tasks. We compare the performance with and without Sketchpad, as well as other notable multimodal LMs, including Gemini [42], Claude 3 [2], and the open-source LLaVA 1.5 [26], LLaVA-NeXT [27].
Model $V^{*}$ench MMVP Depth Spatial Jigsaw Vis. Corr. Sem. Corr.
Prior multimodal LLMs
LLaVA-1.5-7B [26] 48.7 - 52.4 61.5 11.3 25.6 23.0
LLaVA-1.5-13B [26] - 24.7 53.2 67.8 58.0 29.1 32.4
LLaVA-NeXT-34B [27] - - 67.7 74.8 54.7 30.8 23.7
Claude 3 OPUS [2] - - 47.6 58.0 32.7 36.6 25.2
Gemini-Pro [42] 48.2 40.7 40.3 74.8 57.3 42.4 26.6
GPT-4V-preview [35] 55.0 38.7 59.7 72.7 70.0 33.7 28.8
Previous state of the art 75.4 [51] 49.3 [10] 67.7 [27] 76.2 [43] 70.0 [35] 42.4 [42] 33.1 [49]
Latest multimodal LLMs + Visual Sketchpad
GPT-4 Turbo 52.5 71.0 66.1 68.5 64.7 48.8 30.9
+ Sketchpad 71.0 73.3 68.5 80.4 68.5 52.3 42.4
+18.5 +2.3 +2.4 +11.9 +3.8 +3.5 +11.5
GPT-4o 66.0 85.3 71.8 72.0 64.0 73.3 48.6
+ Sketchpad 80.3 86.3 83.9 81.1 70.7 80.8 58.3
+14.3 +1.0 +12.1 +9.1 +6.7 +7.5 +9.7
Table 2: Accuracy on complex visual reasoning tasks. Sketchpad enhances both GPT-4 Turbo and GPT-4o performance, establishing new SOTA performance levels on all the tasks.
Main results. Table 2 shows the performance of our Sketchpad and baselines. Sketchpad consistently improves base model performance across all tasks. GPT-4o with Sketchpad sets the new state-of-the-art results on all tasks. Sketchpad is particularly effective on $V^{*}$ench, yielding 18.5% accuracy improvement for GPT-4 Turbo and 14.3% improvement for GPT-4o, surpassing the previous state of the art SEAL [51] which used a visual search model specifically trained for this task. On BLINK tasks, Sketchpad on average yields 6.6% absolute accuracy gain for GPT-4 Turbo and 9.0% gain for GPT-4o. Interestingly, despite the fact that all modules in Sketchpad work on a single image, the LMs also get substantial improvement on multi-image tasks, including jigsaw puzzles, visual correspondence, and semantic correspondence. Finally, GPT-4o, the LM with stronger multimodal ability than GPT-4 Turbo, benefits more from Sketchpad. For example, on the relative depth task, GPT-4o gets 12.1% accuracy improvement, while GPT-4 Turbo only gets 2.4%, showing that GPT-4o is better at understanding the depth map Sketchpad generated. Overall, our experiments show that Sketchpad is an effective way to improve multimodal LMs’ performance on visual reasoning tasks.
How many times is each vision specialist used? We count the number of times each vision specialist is used in each task, as shown in Figure 3. Here we choose the four tasks that achieve the largest improvement: $V^{*}$ench, relative depth, spatial reasoning, and semantic correspondence. We observe that (1) the use of vision specialist is task-dependent, and the two LMs analyzed utilize similar tools. For example, for $V^{*}$, which needs to locate small objects, the LMs mainly use detection, sliding window search, and zoom-in, similar to how people would search. For the relative depth task, both models rely on depth estimation. For spatial reasoning, the LMs use detection and segmentation to facilitate visual reasoning. (2) GPT-4o likes to use more tools. GPT-4o uses the vision specialists more often than GPT-4 Turbo. Also, the two LMs behave differently for the semantic correspondence tasks. GPT-4o uses the segmentation module for $40\%$ of the task instances, while GPT-4 Turbo uses the detection module for less than $20\%$ of times, and rarely uses the segmentation module. This difference may explain the performance gap between the two LMs (58.3% v.s. 42.4%) on this task.
[…]
figurePercentage of times GPT-4o and GPT-4 Turbo use a visual module in Sketchpad when solving $V^{*}$ench, relative depth, spatial reasoning, and semantic correspondence tasks.
Model $V^{*}$ MMVP Depth Spatial
GPT-4 Turbo 52.5 71.0 66.1 68.5
    SoM 42.0 60.7 58.9 78.3
    SoM + orig. 51.3 74.3 66.9 79.7
    Visprog 33.2 16.3 67.8 53.8
    Sketchpad 71.0 73.3 68.5 80.4
GPT-4o 66.0 85.3 71.8 72.0
    SoM 49.0 70.7 62.9 83.2
    SoM + orig. 68.1 84.0 75.0 82.5
    Visprog 32.4 17.3 46.8 37.8
    Sketchpad 80.3 86.3 83.9 81.1
Table 3: Comparison with other augmentation frameworks for multimodal LMs on single-image tasks. For fair comparison, we modify the original Visprog [14] framework by replacing the LM and VQA components with the corresponding GPT-4 model.
Comparison with visual prompting and tool-use frameworks. In Table 3, we compare Sketchpad with the visual prompting framework SoM [55] and the LLM tool-use framework Visprog [14]. Details of these methods can be found in §2. For a fair comparison, we make the following adaptations: (1) we find that prompting LMs with SoM images can hurt performance, likely because the visual prompts confuse the model. To make a stronger baseline, we prompt the LM with both the original image and the SoM image (full prompt in §C), which we refer as “SoM + orig.” (2) We replace the LM and VQA modules in Visprog with the corresponding GPT-4 model. (3) Since baseline methods are developed on single-image tasks, we compare Sketchpad on such tasks. From Table 3, we can see that Sketchpad is the only framework that yields consistent improvement on all tasks. SoM can boost spatial reasoning ability, as the authors reported. However, it can hurt the performance on other tasks, even in the “SoM + orig.” setting. Visprog performs worse than the base LM on all the tasks. As prior work [21, 18] suggests, one possible reason is that the vision modules themselves have errors, and the error propagates when the modules are composed by a program.
6 Analysis and Discussion
Why does Sketchpad work? First, vision is a versatile and informational interface that complements language. Dense information like depth and segmentation cannot be described easily through language [9]. In a broader perspective, humans have developed many visualization techniques that are direct, efficient, and informational. Sketchpad provides LMs the opportunity to use them. Second, in Sketchpad, multimodal LMs can plan and reason based on the intermediate visual artifacts they created. In contrast, in prior modular vision work [14, 38, 55], multimodal modules follow a predefined plan by either humans or code. Sketchpad is much more flexible and robust to errors. For example, suppose object detection makes an error. The LM can (in principle) find the error by viewing the bounding boxes, and change its following plans, but prior methods cannot. Third, as discussed next, the plans of multimodal LMs are similar to human plans, and therefore likely benefit from the fact that the underlying LMs have seen data with similar reasoning patterns.
Do LMs have the same plans as humans? We conduct a human study on all geometry problems and 10 problems on each vision task. On geometry, humans draw the same auxiliary line as GPT-4o 80% of the time. On vision, we show 2 human subjects the full plan of GPT-4o, which they rate is valid in 92.8% of instances. Most errors are caused by failures in the vision specialists (e.g., fail to detect an object) and mistakes in simple visual question answering, rather than planning.
Model Geometry Maxflow Convexity Winner ID
LLaVA-NeXT-13B 11.1 7.8 50.39 5.8
+ oracle Sketchpad 22.2 10.2 50.0 36.7
LLaVA-NeXT-34B 26.1 0.8 81.6 49.0
+ oracle Sketchpad 28.3 14.1 87.1 49.4
Table 4: Open-source LLaVA models’ performance on math tasks. The oracle Sketchpad uses the visual artifact generated in the last action of GPT-4o + Sketchpad as inputs.
Experiments on open-source models. Can sketches like diagrams, plots, and auxiliary lines facilitate existing open-source multimodal LMs? To answer this question, we conduct the experiments in Table 4. We use the visual artifacts generated in the last action of GPT-4o + Sketchpad experiment as the image input for open-source LLaVA-NEXT models [27]. We can see that this oracle Sketchpad brings consistent improvement to math tasks and boosts mathematical reasoning.
7 Conclusion
We present Visual Sketchpad, a framework that provides multimodal LMs with the tools necessary to generate intermediate sketches to reason over tasks. For complex mathematical reasoning tasks, Sketchpad yields large performance gains, by visualizing auxiliary lines, math functions, graphs, and games during reasoning. For visual reasoning tasks, we add vision specialists to Sketchpad. The LM can call these specialists during reasoning, observing the visualization of these specialists’ predictions (e.g., bounding boxes from the object detection model; masks from the segmentation model), and then conduct further planning and reasoning. Experiments show that Sketchpad enhances the LMs’ performance across all tasks, and sets new state-of-the-art results. Ultimately, Sketchpad represents a step toward endowing LMs with more human-like multimodal intelligence, leveraging the complementary strengths of language and vision to tackle increasingly complex reasoning challenges.
Limitations and future directions. First, Sketchpad requires more computing resources than directly outputting language tokens. We discuss more about computing costs in E. Second, this work focuses on existing off-the-shelf LMs. Future work may explore the training side of Sketchpad. For example, recent multimodal models like Unified-IO 2 [31] and Chameleon [41] are natively multimodal and can output both text and images. Sketchpad may emerge as a new paradigm for instruction tuning these models. Finally, Sketchpad can be applied in more areas. For example, for robotics, we can apply Sketchpad to search for small things in a crowded space, highlight the object of interest, and zoom the camera for a better view or use depth estimation to help navigation.
Appendix A Acknowledgement
This work is in part supported by gifts from Allen Institute for Artificial Intelligence, Meta, Apple, and Amazon Science. We also greatly appreciate the help from Zixian Ma, Wei-Chiu Ma, Benlin Liu, Jieyu Zhang, UW NLP, AllenNLP, AI2 PRIOR, and UW RAIVN for their help.
Appendix B More Qualitative Examples
Here we provide more details of how GPT-4o and Sketchpad solve complex mathematical and visual reasoning tasks.
Figure 4: Qualitative examples of how GPT-4o and Sketchpad solve (a) complex mathematical and (b) visual reasoning tasks.
Appendix C Prompts
Prompts for computer vision tasks. Here we provide the running log of GPT-4o + Sketchpad tackling a visual search problem from $V^{*}$ench [51]. We use the same prompt template for all computer vision problems in this work. For visualization purposes, we present the prompts that contain codes differently. In our implementation, they are all text inputs to the LLM.
SYSTEM PROMPT You are a helpful multimodal AI assistant. [MORE INSTRUCTIONS …]    For each turn, you should first do a "THOUGHT", based on the images and text you see. If you think you get the answer to the intial user request, you can reply with "ANSWER: <your answer>" and ends with "TERMINATE".
Initial Prompt + Request
1Here are some tools that can help you. All are python codes. They are in tools.py and will be imported for you.
2The images has their own coordinate system. The upper left corner of the image is the origin (0, 0). All coordinates are normalized, i.e., the range is [0, 1].
3All bounding boxes are in the format of [x, y, w, h], which is a python list. x is the horizontal coordinate of the upper-left corner of the box, y is the vertical coordinate of that corner, w is the box width, and h is the box height.
4Notice that you, as an AI assistant, is not good at locating things and describe them with coordinate. You can use tools to generate bounding boxes.
5You are also not good at answering questions about small visual details in the image. You can use tools to zoom in on the image to see the details. Below are the tools in tools.py:
6“‘python
7class AnnotatedImage:
8 # A class to represent an annotated image. It contains the annotated image and the original image.
9
10 def __init__(self, annotated_image: Image.Image, original_image: Image.Image=None):
11 self.annotated_image = annotated_image
12 self.original_image = original_image
13
14def detection(image, objects):
15 """Object detection using Grounding DINO model. It returns the annotated image and the bounding boxes of the detected objects.
16 The text can be simple noun, or simple phrase (e.g., ’bus’, ’red car’). Cannot be too hard or the model will break.
17 The detector is not perfect, it may wrongly detect objects or miss some objects.
18 Also, notice that the bounding box label might be out of the image boundary.
19 You should use the output as a reference, not as a ground truth.
20 When answering questions about the image, you should double-check the detected objects.
21
22 Args:
23 image (PIL.Image.Image): the input image
24 objects (List[str]): a list of objects to detect. Each object should be a simple noun or a simple phrase. Should not be hard or abstract concepts like "text" or "number".
25
26 Returns:
27 output_image (AnnotatedImage): the original image, annotated with bounding boxes. Each box is labeled with the detected object, and an index.
28 processed boxes (List): listthe bounding boxes of the detected objects
29
30 Example:
31 image = Image.open("sample_img.jpg")
32 output_image, boxes = detection(image, ["bus"])
33 display(output_image.annotated_image)
34 print(boxes) # [[0.24, 0.21, 0.3, 0.4], [0.6, 0.3, 0.2, 0.3]]
35 # you need to double-check the detected objects. Some objects may be missed or wrongly detected.
36 """
37
38def sliding_window_detection(image, objects):
39 """Use this when you are searching for objects in the image, but the objects are not detected by the object detection model.
40 In that case, the most common reason is that the object is too small such that both the vision-language model and the object detection model fail to detect it.
41 This function tries to detect the object by sliding window search.
42 With the help of the detection model, it tries to detect the object in the zoomed-in patches.
43 The function returns a list of annotated images that may contain at leas one of the objects, annotated with bounding boxes.
44 It also returns a list of a list of bounding boxes of the detected objects.
45
46 Args:
47 image (PIL.Image.Image): the input image
48 objects (List[str]): a list of objects to detect. Each object should be a simple noun or a simple phrase. Should not be hard or abstract concepts like "text" or "number".
49
50 Returns:
51 possible_patches (List[AnnotatedImage]): a list of annotated zoomed-in images that may contain the object, annotated with bounding boxes.
52 possible_boxes (List[List[List[Float]]]): For each image in possible_patches, a list of bounding boxes of the detected objects.
53 The coordinates are w.r.t. each zoomed-in image. The order of the boxes is the same as the order of the images in possible_patches.
54
55 Example:
56 image = Image.open("sample_img.jpg")
57 possible_patches, possible_boxes = search_object_and_zoom(image, ["bird", "sign"])
58 for i, patch in enumerate(possible_patches):
59 print(f"Patch {i}:")
60 display(patch.annotated_image)
61
62 # print the bounding boxes of the detected objects in the first patch
63 print(possible_boxes[0]) # [[0.24, 0.21, 0.3, 0.4], [0.6, 0.3, 0.2, 0.3]]
64 """
65
66def segment_and_mark(image, anno_mode:list = [’Mask’, ’Mark’]):
67 """Use a segmentation model to segment the image, and add colorful masks on the segmented objects. Each segment is also labeled with a number.
68 The annotated image is returned along with the bounding boxes of the segmented objects.
69 This tool may help you to better reason about the relationship between objects, which can be useful for spatial reasoning etc.
70 DO NOT use this tool to search or detect an object. It is likely the object is small and segmentaiton does not help.
71 Segmentation and marking can also be helpful for 3D and video reasoning. For example, helping you to see more clearly and analyzes the relationship between different frames of a video.
72
73 Args:
74 image (PIL.Image.Image): the input image
75 anno_mode (list, optional): What annotation is added on the input image. Mask is the colorful masks. And mark is the number labels. Defaults to [’Mask’, ’Mark’].
76
77 Returns:
78 output_image (AnnotatedImage): the original image annotated with colorful masks and number labels. Each mask is labeled with a number. The number label starts at 1.
79 bboxes (List): listthe bounding boxes of the masks.The order of the boxes is the same as the order of the number labels.
80
81 Example:
82 User request: I want to find a seat close to windows, where should I sit?
83 Code:
84 “‘python
85 image = Image.open("sample_img.jpg")
86 output_image, bboxes = segment_and_mark(image)
87 display(output_image.annotated_image)
88 “‘
89 Model reply: You can sit on the chair numbered as 5, which is close to the window.
90 User: Give me the bounding box of that chair.
91 Code:
92 “‘python
93 print(bboxes[4]) # [0.24, 0.21, 0.3, 0.4]
94 “‘
95 Model reply: The bounding box of the chair numbered as 5 is [0.24, 0.21, 0.3, 0.4].
96 """
97
98def depth(image):
99 """Depth estimation using DepthAnything model. It returns the depth map of the input image.
100 A colormap is used to represent the depth. It uses Inferno colormap. The closer the object, the warmer the color.
101 This tool may help you to better reason about the spatial relationship, like which object is closer to the camera.
102
103 Args:
104 image (PIL.Image.Image): the input image
105
106 Returns:
107 output_image (PIL.Image.Image): the depth map of the input image
108
109 Example:
110 image = Image.open("sample_img.jpg")
111 output_image = depth(image)
112 display(output_image)
113 """
114
115def zoom_in_image_by_bbox(image, box, padding=0.05):
116 """A simple wrapper function to crop the image based on the bounding box.
117 When you want to answer question about visual details in a bounding box annotated by the detection tool, you would like to zoom in on the object using this function.
118
119 Args:
120 image (PIL.Image.Image): the input image
121 box (List[float]): the bounding box in the format of [x, y, w, h]
122 padding (float, optional): The padding for the image crop, outside of the bounding box. Defaults to 0.1. The zoom factor cannot be too small. Minimum is 0.05
123
124 Returns:
125 cropped_img (PIL.Image.Image): the cropped image
126
127 Example:
128 image = Image.open("sample_img.jpg")
129 annotated_img, boxes = detection(image, "bus")
130 cropped_img = zoom_in_image_by_bbox(image, boxes[0], padding=0.05)
131 display(cropped_img)
132 """
133
134def overlay_images(background_img, overlay_img, alpha=0.3, bounding_box=[0, 0, 1, 1]):
135 """
136 Overlay an image onto another image with transparency.
137 This is particularly useful visualizing heatmap while preserving some info from the original image.
138 For example, you can overlay a segmented image on a heatmap to better understand the spatial relationship between objects.
139 It will also help seeing the labels, circles on the original image that may not be visible on the heatmap.
140
141 Args:
142 background_img_pil (PIL.Image.Image): The background image in PIL format.
143 overlay_img_pil (PIL.Image.Image): The image to overlay in PIL format.
144 alpha (float): Transparency of the overlay image.
145 bounding_box (List[float]): The bounding box of the overlay image. The format is [x, y, w, h]. The coordinates are normalized to the background image. Defaults to [0, 0, 1, 1].
146
147 Returns:
148 PIL.Image.Image: The resulting image after overlay, in PIL format.
149 s
150 Example:
151 image = Image.open(’original.jpg’)
152 depth_map = depth(image)
153 overlayed_image = overlay_images(depth_map, image, alpha=0.3)
154 display(overlayed_image)
155 """
156“‘
157# GOAL #: Based on the above tools, I want you to reason about how to solve the # USER REQUEST # and generate the actions step by step (each action is a python jupyter notebook code block) to solve the request.
158You may need to use the tools above to process the images and make decisions based on the visual outputs of the previous code blocks.
159Your visual ability is not perfect, so you should use these tools to assist you in reasoning about the images.
160
161# [six in-context examples here]
162
163# USER REQUEST #:
1Is the red balloon above or below the white balloon? Options: (A) The red balloon is below the white balloon. (B) The red balloon is above the white balloon.
2
3# USER IMAGE stored in image_1 as PIL image.
4Now please generate only THOUGHT 0 and ACTION 0 in RESULT. If no action needed, also reply with ANSWER: <your answer> and ends with TERMINATE in the RESULT:
5
6# RESULT #:
Thought + Action 0
1THOUGHT 0: To determine the relative positions of the red and white balloons, I need to detect the balloons in the image. I will use the detection tool to find the balloons.
2ACTION 0:
3“‘python
4image_1_balloon_detection, image_1_balloon_boxes = detection(image_1, ["red balloon", "white balloon"])
5display(image_1_balloon_detection.annotated_image)
6“‘
Observation 0      Execution success. The output is as follows.
Please generate the next THOUGHT and ACTION. If you can get the answer, please also reply with ANSWER: <your answer> and ends with TERMINATE.
Thought + Action 1
1THOUGHT 1: The detection tool did not annotate the balloons in the image. I will use the sliding_window_detection tool to search for the red and white balloons in the image.
2
3ACTION 1:
4“‘python
5possible_patches, possible_boxes = sliding_window_detection(image_1, ["red balloon", "white balloon"])
6for i, patch in enumerate(possible_patches):
7 print(f"Patch {i+1}:")
8 display(patch.annotated_image)
9“‘
Observation 1      Execution success. The output is as follows. Patch 1:
Patch 2:
Patch 3:
Patch 4:
Patch 5:
Please generate the next THOUGHT and ACTION. If you can get the answer, please also reply with ANSWER: <your answer> and ends with TERMINATE.
Thought + Action 2      THOUGHT 2: From Patch 4 and 5, I can see both the red and white balloons. The red balloon is clearly below the white balloon in these patches.
ACTION 2: No action needed.
ANSWER: The red balloon is below the white balloon. So the answer is (A). TERMINATE
Prompts for math tasks. The prompts are similar to the vision task for math tasks, except we remove the computer vision specialists, and add example codes for math plotting. Besides, the user query is different. For each task, the user query part of the prompt is as follows.
PROMPT You are given a real-valued, scalar function $f(x)$. YOUR TASK is to determine whether $f(x)$ is an even function, an odd function, or neither. Definition of an odd function: A function such that $f(-x)=-f(x)$ where the sign is reversed but the absolute value remains the same if the sign of the independent variable is reversed. A function is neither even nor odd if it does not satisfy either condition. Here is the expression of $f(x)$: $f(x)=\frac{-2x^{5}}{2x^{8}-4x^{6}+12x^{4}+4x^{2}+11.16}$ Respond with ‘even’, ‘odd’, ‘neither’ first on whether the function $f(x)$ is even, odd, or neither, based on the definitions and your observation of the function. You can generate matplotlib code to visualize the function. If you can get the answer, please reply with ANSWER: <your answer>, extract the final answer in FINAL ANSWER: <final answer> and ends with TERMINATE in the RESULT. Answer:
Figure 5: Prompt for the Math Parity task. We follow a similar prompt format to [8], except prompting the models to write the code to generate images.
PROMPT You are given a real-valued, scalar function $f(x)$. YOUR TASK is to determine whether $f(x)$ is an convex function or concave function. Definition of a convex function: A function such that $\forall x,y,0\leq t\leq 1,f(tx+(1-t)y)\leq tf(x)+(1-t)f(y)$ Definition of a concave function: A function such that $\forall x,y,0\leq t\leq 1,f(tx+(1-t)y)\geq tf(x)+(1-t)f(y)$ Here is the expression of $f(x)$: $f(x)=7.57-0.08*Abs(x)$ Respond with ‘convex’ or ‘concave’ first on whether the function f (x) is convex or concave, based on the definitions and your observation of the function. You can generate matplotlib code to visualize the function. If you can get the answer, please reply with ANSWER: <your answer>, extract the final answer in FINAL ANSWER: <final answer> and ends with TERMINATE in the RESULT. Answer:
Figure 6: Prompt for the Math Convexity task. We follow the similar prompt format to [8], except prompting the models to write the code to generate images.
PROMPT You are given an adjacency matrix of a graph and two query nodes. YOUR TASK is to find if there is a path between the two nodes. Definition of connectivity: In an undirected graph G, two vertices u and v are called connected if G contains a path from u to v. A path in a graph is a finite sequence of edges which joins a sequence of vertices. In the query example, the nodes and the adjacency matrix are zero-indexed. Query Example: Adjacency Matrix: $\begin{bmatrix}0&0&0&0&0&1&0&0&0\\ 0&0&1&0&0&0&1&0&0\\ 0&1&0&0&1&0&0&0&0\\ 0&0&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 1&0&0&0&0&0&1&0&1\\ 0&1&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\end{bmatrix}$ Query nodes indices (zero-indexed): 4 and 0 Respond with ‘yes’ or ‘no’ first on whether the query nodes are connected or not in the graph. If there is a path, first provide the path as a sequence of vertices (nodes), and then explain your reasoning. You can use networkx to draw the graph. If there is no path, explain why in details. Answer (start with ’yes’ or ’no’): If you can get the answer, please reply with ANSWER: <your answer>, extract the final answer in FINAL ANSWER: <final answer> and ends with TERMINATE in the RESULT. Answer:
Figure 7: Prompt for the Graph Connectivity task. We follow the similar prompt format to [8], except prompting the models to write the code to generate images.
PROMPT You are given a visual representation of two graphs, graph G on the left and graph H on the right. YOUR TASK is to determine whether the two graphs are isomorphic to each other. Definition of graph isomorphism: In graph theory, an isomorphism of graphs G and H is a bijection $f$ between the vertex sets of G and H, denoted as $f:V(G)\rightarrow V(H)$. G and H are said to be isomorphic when $f$ satisfies the following: any two vertices $u$ and $v$ of G are adjacent in G if and only if $f(u)$ and $f(v)$ are adjacent in H. This kind of bijection is commonly described as "edge-preserving bijection", in accordance with the general notion of isomorphism being a structure-preserving bijection. In the query example, the adjacency matrices are zero-indexed. Adjacency Matrix G: $\begin{bmatrix}0&0&0&0&0&1&0&0&0\\ 0&0&1&0&0&0&1&0&0\\ 0&1&0&0&1&0&0&0&0\\ 0&0&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 1&0&0&0&0&0&1&0&1\\ 0&1&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\end{bmatrix}$ Adjacency Matrix H: $\begin{bmatrix}0&0&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&0\\ 1&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&1&0&0\\ 0&1&0&1&0&1&0&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&1&0&0&0\end{bmatrix}$ Respond with ’yes’ or ’no’ first on whether the two graphs are isomorphic to each other. You can use networkx to draw the graph. If they are isomorphic, first provide the bijection between the two graphs, and then explain your reasoning. You can use networkx to draw the graph. If they are not isomorphic, explain why in detail. Answer (start with ’yes’ or ’no’): If you can get the answer, please reply with ANSWER: <your answer>, extract the final answer in FINAL ANSWER: <final answer> and ends with TERMINATE in the RESULT. Answer:
Figure 8: Prompt for the Graph Isomorphism task. We follow a similar prompt format to [8], except prompting the models to write the code to generate images.
PROMPT You are given an adjacency matrix of a graph and two query nodes (one source node and one sink node). The source node is the node where the flow starts and the sink node is the node where the flow ends. YOUR TASK is to solve the maxflow problem given the weighted directed graph. Definition of Maxflow problem: In the max flow problem, we have a directed graph with a source node $s$ and a sink node $t$, and each edge has a capacity (integer valued, colored in green) that represents the maximum amount of flow that can be sent through it. The goal is to find the maximum amount of flow that can be sent from $s$ to $t$, while respecting the capacity constraints on the edges. Query Example: Adjacency Matrix: $\begin{bmatrix}0&1&4\\ 0&0&6\\ 0&0&0\end{bmatrix}$ Source node (zero-indexed): 0 Sink node (zero-indexed): 2 In the query example, the nodes and the adjacency matrix are zero-indexed. You can use networkx to draw the graph. If you can get the answer, please reply with ANSWER: <your answer>, extract the final answer in FINAL ANSWER: <final answer> and ends with TERMINATE in the RESULT. Answer:
Figure 9: Prompt for Graph Maxflow task. We follow the similar prompt format to [8], except prompting the models to solve the maxflow problem.
PROMPT Given the following FEN of the chess game: 1r1q1rk1/1b2b1Qp/4pp1B/pp1nP3/2pPN3/P1P5/1PB3PP/R4RK1 b - - 0 18 Determine the game’s outcome. Who won: White or Black? Answer (start with ’white’ or ’black’ or ’draw’):You can draw the chess board using Python given the FEN string. If you can get the answer, please reply with ANSWER: <your answer>, extract the final answer in FINAL ANSWER: <final answer> and ends with TERMINATE in the RESULT. Answer:
Figure 10: Prompt for Winner ID task. We follow a similar prompt format to [8], except prompting the models to analyze the game outcome.
Appendix D Dataset Statistics
Table 5 and  6 show the statistics of the datasets we used, including IsoBench [8], BLINK [9], MMVP [44], and $V^{*}$ench [51].
Dataset size partition representation
Math Parity 383 val code
Math Convexity 255 val code
Graph Maxflow 128 val array
Graph Connectivity 128 val array
Graph Isomorphism 128 val array
Winner ID 257 val FEN
Table 5: IsoBench [8] data statistics.
Dataset size partition input
$V*$Bench 257 - Single Image
MMVP 300 - Single Image
BLINK Relative Depth 124 val Single Image
BLINK Spatial Relation 143 val Single Image
BLINK Jigsaw Puzzle 150 val Multiple Images
BLINK Visual Correspondence 172 val Multiple Images
BLINK Semantic Correspondence 139 val Multiple Image
Table 6: Vision tasks data statistics.
Appendix E Costs
The cost of running each task using GPT-4o is in Table 7.
Dataset tokens per sample GPT-4o cost per sample
Math Parity 2994 \$0.015
Math Convexity 2211 \$0.011
Graph Connectivity 2819 \$0.014
Graph Isomorphism 3143 \$0.016
$V^{*}$ench 26647 \$0.133
MMVP 11870 \$0.059
BLINK Relative Detph 14078 \$0.070
BLINK Spatial Relation 12848 \$0.064
BLINK Jigsaw Puzzle 13206 \$0.066
BLINK Visual Correspondence 16988 \$0.085
BLINK Semantic Correspondence 11508 \$0.058
Table 7: The cost of running Sketchpad on each task.
Appendix F Impact Statement
Our work proposes Sketchpad, a framework aiming at advancing academic research and meeting industry needs. In a broader perspective, Sketchpad proposes a new way that humans can interact with LMs, and makes LMs more interpretable by eliciting their reasoning process with both language and sketches. On the other hand, if misused, the LMs may be used to generate harmful vision and text artifacts. Nevertheless, this is not directly related to our research, and more researchers can be involved to research on the safety issue in a multimodal context.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes] .
  4. Justification: Both the abstract and introduction reflect our paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of this work in section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not make any theoretical claims.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide implementation details in §3 and §C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The dataset is accessible online. We are committed to releasing the code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The test details are specified in D, 4 and 5.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Since our work focus on inference of language models, the results are provided as evaluation scores on validation/test set. And we have explained how they are calculated in 4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details of computing resources like compute device, number of workers, memory and time of execution in §E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics, and strictly followed it in this work.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, we talk about boarder impacts in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper focus on inference of language models, and thus there is no data or models to release.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes. We acknowledge the codebases we have referred to, cite and provide the urls of datasets and models in LABEL:sec:implementations.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.